OpenStudy (anonymous):

all edges of a cube are expanding at a rate of 3 centimeters per second. how fast is the volume changing when each edge is (a) 1 cm. (b) 10 cm.

OpenStudy (anonymous):

looks like a related rate problem fortunately the formula for the volume of a cube is easy, it is \(V(x)=x^3\)

OpenStudy (anonymous):

take the derivative with respect to time and get \[V'=3x^2x'\]

OpenStudy (anonymous):

you are told \(x'=3\) so this is really \[V'=9x^2\] replace \(x\) by the various values to get \(V'\)

OpenStudy (anonymous):

dV/dt=9x^2dx/dt ?:)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i used \(V'\) and \(x'\) instead of \(\frac{dV}{dt}\) and \(\frac{dx}{dt}\) because it was easier for me to write also for you it is easier

OpenStudy (anonymous):

but in your notation, you have \[V=x^3\] so \[\frac{dV}{dt}=3x^2\frac{dx}{dt}\]

OpenStudy (anonymous):

this line all edges of a cube are expanding at a rate of 3 centimeters per second tells you that \(\frac{dx}{dt}=3\)

OpenStudy (anonymous):

that means \(\frac{dV}{dt}=3x^2\times 3=9x^2\)

OpenStudy (anonymous):

let me know if i lost you there

OpenStudy (anonymous):

Ok got it(:

OpenStudy (anonymous):

So dats da answer?:)